y = -2x + 19
y = x + 7

Answer by stanbon(75887)

y = -2x + 19
y = x + 7
-------
Substitute and solve for "x"::
-2x+19 = x+7
3x = 12
x = 4
----
Solve for "y":
y = x + 7
y = 4+7 = 11
------------------
Cheers,
Stan H.
----------------